# Privacy in Speech Technology

### v3

@@ III-D Attack and attacker properties
We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences. A malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a honest-but-curious attacker seizes emerging opportunities [99, 100]. The attack surface or vector is the specific point in the system where private information leaves the authorized path to make its way to the attacker [101].
@@ IV-A Vulnerable data sources
Speech data can be exposed in two ways (see fig. 2): First, when a target information is transmitted, it is a threat to privacy when contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party. Second, since speech contains a wide range of private information (see table I) which is bundled in the acoustic signal in complicated ways, it is challenging to extract only the desired message. Typically, speech always contains consequential side information bundled into the target information. It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the target information. Per definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
The difference between the two forms of information makes an important difference in approaches to mitigation. Transmitting the target information to the recipient is the desired outcome of the interaction and we can thus not complete the desired task without sending the message. In some cases, though, we can reduce the amount of information content in the target information, and more generally, we should always secure the transmission for example through cryptography (see sections V-A, V-B and V-F). Some cryptographic methods, such as homomorphic encryption, even keep the message secret from the receiver but allow the processing of a desired task. However, when side-information is exposed, as additional methods we can also use signal processing to better remove, replace, or distort such side-information (see section V-A).
@@ IV-B Attack types
There are three main types of attacks; motivated, opportunistic, and incidental. In motivated attacks, the attacker attempts to extract information about a specific person. First, the attacker can try to extract private information from speech samples from that person. For example, an attacker could extract private health information from a public speech of a famous person. Second, an attacker can browse through a database of anonymous speakers in an attempt to identify the target person. This is known as re-identification, as the attacker thus recovers the identity of a speaker from a database where the identities were not available. When the target speaker is re-identified, the attacker can then extract private information from the re-identified sample. Moreover, the fact that the voice of the target person is in the database can be private or sensitive information in itself. For example, if a database contains speech samples from the members of a particular association, then the re-identification of a target person reveals that person’s affiliation with that association.
Opportunistic attacks are usually known as birthday attacks, referring to the birthday paradox [102], where the attacker uses a known weakness on a large pool of potential victims. It is like spam email; if the attacker sends a large enough number of such emails, then randomly, someone triggers the trap by mistake. In this form of attack, anyone could be the victim. For example, a speech service could analyze private speaker attributes (for, e.g., health problems) or affiliations (with, e.g., minority groups), for purposes of targeted advertisement or surveillance. Where motivated attacks typically target high-value victims such as famous persons, where both protections and rewards are high, birthday attacks typically target ordinary people, where the likelihood of success is larger but rewards are smaller.
@@ IV-C Attack model
Figure 3 illustrates an attack model for measuring the extent of privacy [56]. We assume that there is some private data, which is anonymized such that it can be used in a trusted task. The attacker has access to the anonymized data, which is here called “public data”. Observe that the public data is not necessarily openly available for anyone to see, but it only indicates that the data is sufficiently freely available that the attacker has access to it. The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information. Again, the term “found data” is a loose term and denotes any data other than found data that is available to the attacker.
To complement an Attack(er) model, it can also be useful to define a Protector model for the use-case of the system. Jointly, the two can be documented using the Scenario-of-Use scheme, as illustrated in table III [103]. The accurate definition of the protector model and use-cases is essential, since they define both the target information needed as well as how performance should be evaluated [104].
TABLE III: The Scenario-of-Use scheme [103].
@@
Attacks can be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used. Access to speech samples from the target user (known as enrollment data) are particularly important for the attacker. For example, in the anonymization task of the VoicePrivacy 2022 challenge, the objective was to replace (pseudonymize) speaker identity but retain all other speech characteristics such as linguistic information. The anonymized sentences are known as trial utterances. The attack scenarios were then classified as [98]:
@@
    Lazy-informed: Both trial and enrollment data are available to attackers, but anonymized with different pseudo-speaker.
@@
    Semi-informed: As in lazy-informed, attackers can also train their model with the same anonymization system but different pseudo-speakers.
With increasing information, obviously, the attacker’s task becomes easier and the accuracy of the attacker’s models improves. Further, such attack scenarios can then be devised according to the specific use case. For example, if the task is to anonymize emotional or health status, we can define different attack scenarios depending on the extent to which the attacker has information about categories of emotions and health status used in anonymization.
@@
Figure 3: An attack model for the evaluation of privacy-preserving anonymization, where private data is anonymized to remove private information, and the anonymized data is shared publicly. An attacker uses any available (found) data and anonymized public data to infer private information contrary to the users’ preferences. Anonymized data flow is indicated by dashed lines and the attack by red lines and an exclamation mark.
@@ IV-D1 Cloud leak
Another example is data collection for training of deep neural networks. A majority of advanced speech processing today uses machine learning, which has to be trained using large databases of speech. The best quality data correspond closely to scenarios where the services are used. Recording users’ interactions with their devices is then attractive for training improved models since it corresponds exactly to the use case. This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device but potentially also any and all speech in the vicinity of the device.
@@ IV-D2 False activation
Devices with speech interfaces can hear all conversations in the same acoustic space as the device and, therefore, need mechanisms to determine which utterances are intended for the speech interface (see fig. 5). A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [105]. The wake word is then like a rudimentary password, which prevents the interface from activating when speech is not directed to the device. Unfortunately, designing wake word detectors is non-trivial, and they will occasionally make mistakes. They might sometimes miss a wake word when it is spoken (false negative) or mistake some other unrelated sound as a wake word (false positive). While false negatives are annoying for the user when the service does not activate, false positives potentially present serious threats to privacy. In some famous cases, speech interfaces have activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [106, 107].
@@ IV-D3 Cloud access
An example of this threat is medical records, which can be collected from patients to a centralized database. A researcher with authorization to access the database can then potentially extract private information beyond the expected [108].
@@ IV-D4 WASN authentication
The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [109, 79, 110]. Such collaboration can be realized with wireless acoustic sensor networks (WASNs), where several independent devices simultaneously pick up speech, and those channels are combined to obtain a high-quality signal (see fig. 7[16, 18]. A sensitive question is, however, authorization: Which devices are allowed to share information with each other?
One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [18, 72]. Presence in the same space is thus already an implicit authorization to participate in a joint signal pickup. To protect privacy, it is then necessary to determine which devices reside in the same acoustic space. Devices in a different room can belong to the same company or family and be connected to the same network, but they are still outside the sphere of the current discussion. Without proper authorization mechanisms, devices outside the room could then gain access to private speech inside the room.
@@ IV-D5 Speech interface leak
In the case where Eve overhears Alice’s speech, we assume that people have a learned awareness of which other people are present in the same room and adjust their speech accordingly. Then either Alice does not mind that Eve overhears her speech (inconsequential information), Alice can change her speech style to a whisper, or change content such that Eve does not hear anything private (reduced information transmission), or she can go to a different room to continue the interaction in private (modified acoustic channel). Whether users are similarly aware of devices present in the room, through anthropomorphization of devices or otherwise, and adjust their speech accordingly is an open question.
@@
Figure 9: Threat scenario “Discussion leak”, where a dialogue between two users, Alice and Bob, is picked up by a local device and potentially shared to a remote service contrary to preferences (red arrow and exclamation mark).
@@ IV-D6 Discussion leak
Figure 9 illustrates a discussion between two persons overheard by a local device and, like False activation, triggers unintended functionalities in the local device. The main difference to a Speech interface leak (sections IV-D5 and 8) is that the type of speech is different. Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues, whereas spoken human-to-human interaction has potentially a much more dynamic character. This is potentially beneficial in the sense that the speaking style of human-to-human interaction could be recognized, preventing unintended triggering of functionalities.
@@ IV-D7 Shared device
Many practical scenarios include multiple users sharing one or several speech-operated devices. For example, a family can share a smart speaker or television, an office meeting room can have smart devices and customer service points can have a phone shared across duty officers. Each of these devices can collect private information about the user(s) over time and can potentially share it with other users (see fig. 10). Notably, this scenario highlights that leak happens over a distance in time, whereas threats occurring over a spatial distance often receive most of the attention.
@@ V-A Information isolation
By isolation of information, we refer to the design of a channel of information where the target information is isolated from other (undesired) private information. This applies at least and in particular to the attack vectors Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), where a reduction in flow of private information would reduce threats to privacy. Central concepts in information isolation include 1. information bottlenecks(or privacy funnels) , where information is compressed to a low-capacity channel such that only target information is passed through (see section V-A1) and where the neural network implementation usually takes the structure of an autoencoder [111], 2. adversarial training, where an information channel (such as an information bottleneck) is trained to contain target information to fulfill the trusted task and simultaneously trained such that a competing, attacking adversarial model fails to extract private information (threat task), (see section V-A2) and 3. disentanglement, where different attributes of speech are isolated to separate channels, as a generalization of information bottlenecks (see section V-A3). Note that these concepts are overlapping and should be seen as characterizations of algorithms and methodology, rather than alternative approaches.
@@ V-A1 Information bottleneck
The purpose of an information bottleneck is to use an encoder to distill the information needed to perform a downstream (trusted) task into a bottleneck, such that all other (private) information is removed. A decoder can then be used to decipher the information in the bottleneck into a format usable for the downstream task. The overall approach is generic and can be applied, for example, in cases where 1. the decoder resynthesizes the signal with some attributes anonymized [112] or 2. information extraction scenarios such as classification [113], where the decoder can often be trivial.
@@
A concept related to information bottlenecks is the privacy funnel [114]. Where the original definition of the information bottleneck [113] aims to maximize the utility of a channel with a fixed rate, the privacy funnel aims to minimize the information rate with a fixed utility. The information bottleneck is, however, much more widely known in the community and often informally used as an umbrella term to cover both cases.
@@
To distill only the target information and discard any side-information, speech information can be passed through a bottleneck so tight, that only the legitimate message at full accuracy can pass through [115, 114] (see fig. 11). This approach thus provides protection against an attacker who has access to the output of the bottleneck. The challenges are to design a training methodology and model structure with a sufficiently tight bottleneck so that the legitimate message is accurately retained while private side information is discarded. Observe that a generic implementation of an information bottleneck does not guarantee privacy as side-information can sneak through the bottleneck alongside the target information [116]. Instead, to be privacy preserving, a channel must be designed or optimized for privacy, warranting the term privacy funnel.
@@
A frequently used approach to implementing a bottleneck is an autoencoder structure [111], consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [78, 49, 117] (see fig. 12). To be privacy-preserving, the bottleneck should be sufficiently tight that the original speech signal cannot be perfectly reconstructed to resemble the original in its waveform with respect to some distance function. Designing a loss function is then challenging when the bottleneck is so tight that the decoded signal is very far from the original. One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [118, 119].
@@
Information content at the bottleneck can be reduced by either reducing the signal rank (viz. dimensionality reduction, e.g., [49]) or by quantizing and coding the signal at a low bitrate (e.g., [112, 120]). The trade-off between the accuracy of trusted task and bottleneck entropy then determines the extent of privacy (see section VI). Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information to that channel. Since typical representations have $B=32$ or $B=64$, which can encode up to $2^{B}$ unique speaker identities, we see that the bottleneck is not particularly tight. As a general rule, therefore, bottlenecks should always be quantized to the lowest bitrate sufficient to represent the target information.
@@
A central advantage of the information bottleneck approach is that it can give a provable level of privacy. If the bitrate of the bottleneck is $B_{b}$ and that of the target information is $B_{t}$, then the privacy leak is at most $B_{leak}\leq B_{t}-B_{b}$. This applies to single vectors of target information passed through a bottleneck with full accuracy. Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [121, 122]. This approach also addresses only the average disclosure of private information but does not give guarantees of worst-case disclosure [87, 123].
@@
Information bottlenecks, or more accurately, privacy funnels, are thus theoretically well-suited for privacy-preserving applications. They have also been used in other areas, such as privacy-aware transmission of images [124] and face recognition [125]. However, the above-mentioned applications in speech either do not specifically apply it for privacy or use information bottlenecks optimized for privacy with additional components.

### v4

@@ III-D Attack and attacker properties
We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences. A malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a honest-but-curious attacker seizes emerging opportunities [CASEY20038, le2023privacy]. The attack surface or vector is the specific point in the system where private information leaves the authorized path to make its way to the attacker [THEISEN201894].
@@ IV-A Vulnerable data sources
Speech data can be exposed in two ways (see fig. 2): First, when target information is transmitted, it is a threat to privacy when, contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party. Second, since speech contains a wide range of private information (see table I) which is bundled in the acoustic signal in complicated ways, it is challenging to extract only the desired message. Typically, speech always contains consequential side information bundled into the target information. It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the target information. By definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
The difference between the two forms of information makes an important difference in approaches to mitigation. Transmitting the target information to the recipient is the desired outcome of the interaction, and we can thus not complete the desired task without sending the message. In some cases, though, we can reduce the amount of information content in the target information, and more generally, we should always secure the transmission, for example, through cryptography (see sections V-A, V-B and V-F). Some cryptographic methods, such as homomorphic encryption, even keep the message secret from the receiver but allow the processing of a desired task. However, when side-information is exposed, as additional methods we can also use signal processing to better remove, replace, or distort such side-information (see section V-A).
@@ IV-B Attack types
There are three main types of attacks: motivated, opportunistic, and incidental. In motivated attacks, the attacker attempts to extract information about a specific person. First, the attacker can try to extract private information from speech samples from that person. For example, an attacker could extract private health information from a public speech of a famous person. Second, an attacker can browse through a database of anonymous speakers in an attempt to identify the target person. This is known as re-identification, as the attacker thus recovers the identity of a speaker from a database where the identities were not available. When the target speaker is re-identified, the attacker can then extract private information from the re-identified sample. Moreover, the fact that the voice of the target person is in the database can be private or sensitive information in itself. For example, if a database contains speech samples from the members of a particular association, then the re-identification of a target person reveals that person’s affiliation with that association.
In opportunistic attacks, the attacker uses a known weakness on a large pool of potential victims. It is similar to spam email; if the attacker sends a large enough number of such emails, then, randomly, someone triggers the trap by mistake as in a birthday attack [menezes2018handbook]. In this form of attack, anyone could be the victim. For example, a speech service could analyze private speaker attributes (for, e.g., health problems) or affiliations (with, e.g., minority groups), for purposes of targeted advertisement or surveillance. Where motivated attacks typically target high-value victims such as famous persons, where both protections and rewards are high, opportunistic attacks typically target ordinary people, where the likelihood of success is larger but rewards are smaller.
@@ IV-C Attack model
Figure 3 illustrates an attack model for measuring the extent of privacy [maouche20_interspeech]. We assume that there is some private data, which is anonymized such that it can be used in a trusted task. The attacker has access to the anonymized data, which is here called “public data”. Observe that the public data is not necessarily openly available for anyone to see, but it only indicates that the data is sufficiently freely available that the attacker has access to it. The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information. Again, the term “found data” is a loose term and denotes any data other than public data that is available to the attacker.
To complement an Attack(er) model, it can also be helpful to define a Protector model for the use case of the system. Jointly, the two can be documented using the Scenario-of-Use scheme, as illustrated in table III [rahman24_spsc]. The accurate definition of the protector model and use-cases is essential, since they define both the target information needed as well as how performance should be evaluated [meyer2025usecasesvoiceanonymization].
TABLE III: The Scenario-of-Use scheme [rahman24_spsc].
@@
Attacks can be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used. Access to speech samples from the target user (known as enrollment data) is particularly important for the attacker. For example, in the anonymization task of the VoicePrivacy 2022 challenge, the objective was to replace (pseudonymize) speaker identity but retain all other speech characteristics, such as linguistic information. The anonymized sentences are known as trial utterances. The attack scenarios were then classified as [tomashenko2022voiceprivacy]:
@@
    Lazy-informed: Both trial and enrollment data are available to attackers, but anonymized with a different pseudo-speaker.
@@
    Semi-informed: Similar to lazy-informed, attackers can also train their model using the same anonymization system but with different pseudo-speakers.
With increasing information, obviously, the attacker’s task becomes easier, and the accuracy of the attacker’s models improves. Further, such attack scenarios can then be devised according to the specific use case. For example, suppose the task is to anonymize emotional or health status. In that case, we can define different attack scenarios depending on the extent to which the attacker has information about categories of emotions and health status used in anonymization.
@@
Figure 3: An attack model for the evaluation of privacy-preserving anonymization, where private data is anonymized to remove private information, and the anonymized data is shared publicly. An attacker uses any available (found) data and anonymized public data to infer private information contrary to the users’ preferences. Anonymized data flow is indicated by dashed lines, and the attack is represented by red lines with an exclamation mark.
@@ IV-D1 Cloud leak
Another example is data collection for training of deep neural networks. A majority of advanced speech processing today uses machine learning, which has to be trained using large databases of speech. The best quality data correspond closely to scenarios where the services are used. Recording users’ interactions with their devices is then attractive for training improved models since it corresponds precisely to the use case. This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device, but potentially also any and all speech in the vicinity of the device.
@@ IV-D2 False activation
Devices with speech interfaces can hear all conversations in the same acoustic space as the device and, therefore, need mechanisms to determine which utterances are intended for the speech interface (see fig. 5). A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [wang2021wake]. The wake word serves as a rudimentary password, preventing the interface from activating when speech is not directed to the device. Unfortunately, designing wake word detectors is non-trivial, and they will occasionally make mistakes. They might sometimes miss a wake word when it is spoken (false negative) or mistake some other unrelated sound as a wake word (false positive). While false negatives are annoying for the user when the service does not activate, false positives potentially present serious threats to privacy. In some famous cases, speech interfaces have been activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [liptak2017alexa, moscaritolo2018alexa].
@@ IV-D3 Cloud access
An example of this threat is medical records, which can be collected from patients and stored in a centralized database. A researcher with authorization to access the database can then potentially extract private information beyond the expected [Teixeira2018].
@@ IV-D4 WASN authentication
The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [das2021enhancement, zarazaga2020acoustic, sigg2020provable]. Such collaboration can be realized with wireless acoustic sensor networks (WASNs), where several independent devices simultaneously pick up speech, and those channels are combined to obtain a high-quality signal (see fig. 7[Nelus2022, zarazaga2022thesis]. A sensitive question, however, is authorization: Which devices are allowed to share information?
One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [zarazaga2022thesis, koppelmann22_interspeech]. Presence in the same space is thus already an implicit authorization to participate in a joint signal pickup. To protect privacy, it is then necessary to determine which devices reside in the same acoustic space. Devices in a different room can belong to the same company or family and be connected to the same network, but they are still outside the sphere of the current discussion. Without proper authorization mechanisms, devices outside the room could then gain access to private speech inside the room.
@@ IV-D5 Speech interface leak
In the case where Eve overhears Alice’s speech, we assume that people have a learned awareness of which other people are present in the same room and adjust their speech accordingly. Then, either Alice does not mind that Eve overhears her speech (inconsequential information), Alice can change her speech style to a whisper, or change the content such that Eve does not hear anything private (reduced information transmission), or she can go to a different room to continue the interaction in private (modified acoustic channel). Whether users are similarly aware of devices present in the room, through anthropomorphization of devices or otherwise, and adjust their speech accordingly is an open question.
@@
Figure 9: Threat scenario “Discussion leak”, where a dialogue between two users, Alice and Bob, is picked up by a local device and potentially shared with a remote service contrary to preferences (red arrow and exclamation mark).
@@ IV-D6 Discussion leak
Figure 9 illustrates a discussion between two persons overheard by a local device and, like False activation, triggers unintended functionalities in the local device. The main difference to a Speech interface leak (sections IV-D5 and 8) is that the type of speech is different. Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues. In contrast, spoken human-to-human interaction has potentially a much more dynamic character. This is potentially beneficial in that the speaking style of human-to-human interaction can be recognized, thereby preventing unintended triggering of functionalities.
@@ IV-D7 Shared device
Many practical scenarios include multiple users sharing one or several speech-operated devices. For example, a family can share a smart speaker or television, an office meeting room can have smart devices, and customer service points can have a phone shared across duty officers. Each of these devices can collect private information about the user(s) over time and can potentially share it with other users (see fig. 10). Notably, this scenario highlights that a leak happens over a distance in time, whereas threats occurring over a spatial distance often receive most of the attention.
@@ V-A Information isolation
By isolation of information, we refer to the design of a channel of information where the target information is isolated from other (undesired) private information. This applies at least and in particular to the attack vectors Cloud leak (section IV-D1), False activation (section IV-D2), and Cloud access (section IV-D3), where a reduction in flow of private information would reduce threats to privacy. Central concepts in information isolation include 1. information bottlenecks(or privacy funnels) , where information is compressed to a low-capacity channel such that only target information is passed through (see section V-A1) and where the neural network implementation usually takes the structure of an autoencoder [li2023comprehensive], 2. adversarial training, where an information channel (such as an information bottleneck) is trained to contain target information to fulfill the trusted task and simultaneously trained such that a competing, attacking adversarial model fails to extract private information (threat task), (see section V-A2) and 3. disentanglement, where different attributes of speech are isolated to separate channels, as a generalization of information bottlenecks (see section V-A3). Note that these concepts are overlapping and should be seen as characterizations of algorithms and methodology, rather than alternative approaches.
@@ V-A1 Information bottleneck
The purpose of an information bottleneck is to use an encoder to distill the information needed to perform a downstream (trusted) task into a bottleneck, such that all other (private) information is removed. A decoder can then be used to decipher the information in the bottleneck into a format usable for the downstream task. The overall approach is generic and can be applied, for example, in cases where 1. the decoder resynthesizes the signal with some attributes anonymized [wang2019vector] or 2. information extraction scenarios such as classification [tishby2000information], where the decoder can often be trivial.
@@
A concept related to information bottlenecks is the privacy funnel [makhdoumi2014information]. Where the original definition of the information bottleneck [tishby2000information] aims to maximize the utility of a channel with a fixed rate, the privacy funnel seeks to minimize the information rate with a fixed utility. The information bottleneck is, however, much more widely known in the community and often informally used as an umbrella term to cover both cases.
@@
To distill only the target information and discard any side-information, speech information can be passed through a bottleneck so tight that only the legitimate message at full accuracy can pass through [Tishby2015, makhdoumi2014information] (see fig. 11). This approach thus protects against an attacker who has access to the output of the bottleneck. The challenges lie in designing a training methodology and model structure with a sufficiently tight bottleneck, so that the legitimate message is accurately retained while private side information is discarded. Observe that a generic implementation of an information bottleneck does not guarantee privacy as side-information can sneak through the bottleneck alongside the target information [raj2019probing]. Instead, to be privacy preserving, a channel must be designed or optimized for privacy, warranting the term privacy funnel.
@@
A frequently used approach to implementing a bottleneck is an autoencoder structure [li2023comprehensive], consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [malekzadeh2018replacement, perero2022x, lu2013speech] (see fig. 12). To be privacy-preserving, the bottleneck should be sufficiently tight that the original speech signal cannot be perfectly reconstructed to resemble the original in its waveform with respect to some distance function. Designing a loss function is then challenging when the bottleneck is so tight that the decoded signal is very far from the original. One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [neumann2019improving, chorowski2019unsupervised].
@@
Information content at the bottleneck can be reduced by either reducing the signal rank (viz. dimensionality reduction, e.g., [perero2022x]) or by quantizing and coding the signal at a low bitrate (e.g., [wang2019vector, williams2021revisiting]). The trade-off between the accuracy of the trusted task and bottleneck entropy then determines the extent of privacy (see section VI). Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information into that channel. Since typical representations have $B=32$ or $B=64$, which can encode up to $2^{B}$ unique speaker identities, we see that the bottleneck is not particularly tight. As a general rule, therefore, bottlenecks should always be quantized to the lowest bitrate sufficient to represent the target information.
@@
A central advantage of the information bottleneck approach is that it can give a provable level of privacy. If the bitrate of the bottleneck is $B_{b}$ and that of the target information is $B_{t}$, then the privacy leak is at most $B_{leak}\leq B_{t}-B_{b}$. This applies to single vectors of target information passed through a bottleneck with full accuracy. Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [kairouz2015composition, backstrom2025privacydisclosure]. This approach addresses only the average disclosure of private information, without guaranteeing worst-case disclosure [nautsch20_interspeech, vali2024porcupine].
@@
Information bottlenecks, or more accurately, privacy funnels, are thus theoretically well-suited for privacy-preserving applications. They have also been used in other areas, such as privacy-aware transmission of images [sun2024privacyaware] and face recognition [razeghi2024deep]. However, the above-mentioned applications in speech either do not specifically apply it for privacy or use information bottlenecks optimized for privacy with additional components.